# Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Knowledge Distillation and Random Data Erasing

### v1

@@ 3.2.1 Main Results
As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missing scenarios on MIMIC-III dataset. Specifically, it outperforms the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5. Similar trends are observed in other settings. On the other hand, NECHO performs well when medical codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8). This highlights the robust and superior performance of NECHO v2.
This performance gain is attributed to: 1) modifying NECHO for both teacher and student to address varying modality significance, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation and transformer representation random distillation and others, to comprehensively mimic teacher at various representation levels, and 3) simulating random visit information by single point data erasing to minimise data distribution gaps and improve KD. These enables student to imitate the teacher in varied incompleteness, ensuring robust performance effectively.
@@ 3.2.2 Ablation Studies
To evaluate our proposed components, we conduct ablation studies on MIMIC-III data, as detailed in Table 3. We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where medical code representation is either highly dominant or minimal.
We first assess the effectiveness of KD. While $\mathcal{L}_{\text{MWCD}}$ occasionally performs better without it, its consistent use generally enhances performance. The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate supervision. Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial, all of which emphasise systematic KD’s importance to align student’s semantic knowledge to that of the teacher. We also evaluate the efficacy of DA in three scenarios: only distillation, only teacher training, and not at all. Overall performance significantly improves, highlighting the significance of the proposed random single data point erasing under missing visit information.
@@ 3.2.3 Comparative Studies
As shown in Table 4, we compare NECHO v2 with various combinations of teacher and student (original to original, original to modified, adopted MAG) and transformer full distillation under the same settings with the ablation studies. Our proposed methodologies achieves the best overall performance, highlighting the importance of: 1) carefully pairing teacher and student to address varying representation dominance and minimise architectural heterogeneity, and 2) introducing randomness into KD to prevent overfitting and foster effective knowledge transfer.
@@ 4 Conclusion
We tackle uncertain missing sequences for robust multimodal SDP with NECHO v2. Firstly, we modify NECHO to dynamically adjust dominant representations under varying missingness. Next, we design a systematic KD pipeline, adopting the modified NECHO and leveraging multiple distillations including modality-wise contrastive and hierarchical distillation, to comprehensively learn teacher representations from complete data. We also randomly erase single data points per visit to reduce the data distribution discrepancy and enhance KD process. Extensive experiments show the efficacy of NECHO v2 over the existing methodologies. The code will be released upon publication.

### v2

@@ 3.2.1 Main Results
As shown in Table 2, NECHO v2 demonstrates remarkable performance across various missing scenarios on MIMIC-III dataset. Specifically, it outperforms MulT by 0.92%, the original NECHO by 1.52%, its teacher by 3.32%, its student by 1.45%, and UnimodalKD by 1.83% in top-10 accuracy at the balanced missingness of 0.5. Similar trends are observed in other settings. In contrary, NECHO performs well when diagnosis codes are mostly present (0.2) but predicts poorly in scenarios where codes are highly missing (0.8). This highlights the consistent supremacy of our NECHO v2.
This performance gain is attributed to: 1) modifying NECHO for both teacher and student to address varying modality significance, 2) implementing systematic KD, including modality-wise contrastive and hierarchical distillation and transformer representation random distillation, MAG distillation, and dual-level logit distillation, to comprehensively mimic teacher at various representation levels, and 3) simulating random missing visit information by single point data random erasing to minimise data distribution gaps. These enables the student to imitate the teacher in varied incompleteness settings, ensuring considerable performance gain effectively.
@@ 3.2.2 Ablation Studies
To evaluate our proposed components, we conduct ablation studies on MIMIC-III data, as detailed in Table 3. We report two scenarios: a balanced missing ratio of 0.2, and an imbalanced ratios of (0.5, 0.8, 0.2), representing two extremes where diagnosis codes representation is either highly dominant or minimal.
We first assess the effectiveness of KD. Whilst $\mathcal{L}_{\text{MWCD}}$ occasionally performs better without it, its consistent use generally enhances performance. The absence of $\mathcal{L}_{\text{TR2D}}$ and $\mathcal{L}_{\text{MAGD}}$ during distillation significantly deteriorates the performance, highlighting the importance of intermediate supervision. Additionally, $\mathcal{L}_{\text{hrchyLD}}$ is beneficial in performance. The quantitative results validate the importance of all components in systematic KD to align student’s semantic knowledge to that of the teacher.
We also evaluate the efficacy of DA in three scenarios: only distillation, only teacher training, and not at all. Overall performance significantly improves, highlighting the significance of the proposed random single data point erasing under missing visit information. This enhances the teacher’s robustness against missingness during training and minimises data distribution discrepancies during distillation, resulting in a student model that is highly resilient to uncertain data incompleteness.
@@ 3.2.3 Comparative Studies
As shown in Table 4, we compare NECHO v2 with various combinations of teacher and student (original to original, original to modified with MAG) and transformer full distillation under the same settings with the ablation studies. Our proposed methodologies achieves the best overall performance, highlighting the importance of: 1) carefully pairing teacher and student to address varying representation dominance and minimise architectural heterogeneity, and 2) introducing randomness into KD to prevent overfitting and foster effective knowledge transfer.
@@ 4 Conclusion
We tackle uncertain missing sequences for robust multimodal SDP with the proposed NECHO v2. Extensive experiments show the stable superiority of NECHO v2 over the existing methods. The code will be released soon.